OpenStudy (scarlettthecamel):

Solve x^2-5x+6=0 I've forgotten how to solve so some help please?:)

OpenStudy (anonymous):

Do you remember the quadratic formula?

OpenStudy (scarlettthecamel):

umm no lol

jhonyy9 (jhonyy9):

not is necessary because you can factoriz it easy

OpenStudy (scarlettthecamel):

i only can solve when it's like two for example 5x+2=20

jhonyy9 (jhonyy9):

-5=-3+ -2 6=-3*(-2) ok ?

OpenStudy (scarlettthecamel):

but there is 3 exressions

OpenStudy (scarlettthecamel):

ok

jhonyy9 (jhonyy9):

try it hence (x-2)(x-3)=x^2 -5x +6

OpenStudy (scarlettthecamel):

but i need to solve not factorise

jhonyy9 (jhonyy9):

do you can continue it how solve ?

OpenStudy (scarlettthecamel):

idk not rly

jhonyy9 (jhonyy9):

x-2=0 x=2 x-3=0 x=3

jhonyy9 (jhonyy9):

so these mean that 3 and 2 are r

jhonyy9 (jhonyy9):

s of

jhonyy9 (jhonyy9):

are roots of this quadtratic

OpenStudy (scarlettthecamel):

ur confusing me whattt

jhonyy9 (jhonyy9):

hope this is understandably easy

OpenStudy (scarlettthecamel):

idk

OpenStudy (anonymous):

Factoring you get (x-3)(x-2) what values of x would make that zero?

OpenStudy (anonymous):

what minus 3 = 0 and what minus 2 = 0?

jhonyy9 (jhonyy9):

so because this quadratic is factorizing completly easy like (x-2)(x-3) so this mean that not is necessary us quadratic formula to solve it ok ?

OpenStudy (scarlettthecamel):

idk i think i will rely on google atm i don't understand

OpenStudy (anonymous):

think what number minus 3 = 0? and what number minus 2 = 0?

OpenStudy (scarlettthecamel):

idek

OpenStudy (anonymous):

3-3=0 and 2-2 = 0 you answers are 3 and 2

OpenStudy (scarlettthecamel):

oh ok and